Solve the following system of equations. If the system cannot be solved, state that it cannot be solved and explain why.
5x-2y=(-24)
-4y= -48-10x

Answer by checkley77(12844)

5x-2y=(-24)MULTIPLY BY -2 & ADD.
-4y= -48-10x
10X-4Y=-48
-10X+4Y=48
------------------
0+0=0 NO SOLUTION.
THESE 2 EQUATIONS ARE THE SAME. 1 EQUATION WITH 2 UNKNOWNS HAVE NO UNIQUE SOLUTION.